# Word-length curve counting on the once-punctured torus

### v2

@@ III.2   Counting
The exceptional cases are for length $4$ and $5$. For length $4$ there are two cases: $a^{2}b^{2}$ and $aba^{-1}b$, which after renaming of generators give us $8$ conjugacy classes, i.e. |$|\{\gamma\in\mathscr{PC}_{1}(\Sigma_{1,1}\mid i(\gamma)=1,\,\ell_{\omega}(\gamma)=4\}|=8$, and for length $5$ we have only the words of type $ab^{-1}a^{-1}b^{2}$ which give |$|\{\gamma\in\mathscr{PC}_{1}(\Sigma_{1,1})\mid i(\gamma)=1,\,\ell_{\omega}(\gamma)=5\}|=8$.
@@
[…]
@@
Without loss of generality, we assume throughout the proof that $x\leq y$. Let $k\in\mathbb{Z}_{\geq 1}$, and let $[n_{i}]_{i}=[n_{1},\dots,n_{k}]$ be a necklace such that $n_{i}\in\{m,m+1\}$ for all $1\leq i\leq k$, |$|\{i\mid n_{i}=m\}|=x$, and |$|\{i\mid n_{i}=m+1\}|=y$.
First, we prove that if $x\mid y$ then the necklace $[n_{i}]_{i}$ does not have variation $2$. If every run of $m+1$ has size $y/x$, then $[n_{i}]_{i}$ has small variation. If there exist two runs of $m+1$ with sizes differing by at least $3$ (for example, $y/x-1$ and $y/x+2$), then $[n_{i}]_{i}$ does not have variation $2$. Thus, if $[n_{i}]_{i}$ has variation $2$, then the sizes of its runs of $m+1$ can only take values in $\{y/x-1,y/x,y/x+1\}$, as the sum of the runs have to sum $y$ and there are $x$ of them.
Further, if there is one run of size $y/x+1$, then there is at least one run of size $y/x-1$ and there can only be one as otherwise, these give immediately two essential pairs of blocks. Therefore, for a necklace with $2$-variation, there is one run of size $y/x-1$, one of size $y/x+1$, and $x-2$ runs of size $y/x$.
If $x>2$, then the necklace cannot have $2$-variation, as one essential pair of blocks is given straight by the runs with difference $2$, i.e. $\{m,m+1,\overset{y/x-1}{\dots},m+1,m\}$ and $\{m+1,\overset{y/x+1}{\dots},m+1\}$, and another one by extending these sets to an adjacent gap of size $y/x$ (e.g. see Figure 7(a)). When $x=2$, there are only two runs, of sizes $y/x-1$ and $y/x+1$, and one unique possible configuration with these, giving always variation $2$ (see Figure 7(b)).
@@
Moreover, the necklace $[n_{i}]_{i}$ has $2$-variation if and only if the associated run necklace $A[n_{i}]_{i}$ described in Figure 2(c) has. This comes naturally from the map $A$ as two blocks of the form $S_{1}=\{m,n_{j+2},\ldots,n_{j+s-1},m\}$ and $S_{2}=\{m+1,n_{j+2},\ldots,n_{j+s-1},m+1\}$ will map to the same sequence of runs with the first and last one being bigger by one at the second case. Conversely, if there is such a sequence of runs, finding the associated pair of blocks in $[n_{i}]_{i}$ is straightforward.
@@
Now, note that since $S_{1}(0)=1$, $|S_{1}(L)|$ can be uniquely determined by (III.4) and (III.5) for any $L\in\mathbb{Z}_{\geq 1}$. On the other hand, a direct computation shows that the function defined by $L\mapsto\lceil\varphi(L)/2\rceil-1$ maps 00 to $1$, and satisfies (III.4) and (III.5). Hence, for any $L\in\mathbb{Z}_{\geq 1}$, we have
@@
=$=\begin{cases}8\,\varphi(L-4)&\text{for odd }L,\\ 8\left(\varphi(L-4)+1+\left\lceil\varphi(L/2)/2\right\rceil-1\right)&\text{foreven }L.\end{cases}$
@@ IV.   All curves
This arises from the following reasoning. We assume $x_{1}=a$. The exponent $n_{1}$ can be any positive integer, giving a factor $t+t^{2}+\cdots=t/(1-t)$. Next, $n_{2}$ can also be any positive integer, and $x_{2}$ can be chosen between $b$ and $b^{-1}$. This gives a factor $2t/(1-t)$, and so on so forth. Finally, $x_{1}$ can also be $a^{-1}$, $b$, or $b^{-1}$, which gives a factor of $4$.
@@
where the term $1/(1-t)$ corresponds to the case $k=1$. This can be rewritten as
@@
and therefore, as primitive words of length $n$ and primitive words of length $n$ are in $n$-to-$1$ correspondence, the number of primitive reduced necklaces in $\{a,b,a^{-1},b^{-1}\}$ is
@@
$\sum_{d\mid n}\mu(d)=\begin{cases}1&\text{if }n=1,\\ 0&\text{if }n\geq 2,\end{cases}\quad\text{and}\quad\sum_{d\mid n}\mu(d)(-1)^{n/d}=\begin{cases}-1&\text{if }n=1,\\ 2&\text{if }n=2,\\ 0&\text{if }n\geq 3.\\ \end{cases}$

### v3

@@ III.2  Counting
The exceptional cases are for length $4$ and $5$. For length $4$ there are two cases: $a^{2}b^{2}$ and $aba^{-1}b$, which after renaming of generators give us $8$ conjugacy classes, i.e. $|\{\gamma\in\mathscr{PC}_{1}(\Sigma_{1,1}\mid i(\gamma)=1,\,\ell_{\omega}(\gamma)=4\}|=8$, and for length $5$ we have only the words of type $ab^{-1}a^{-1}b^{2}$ which give $|\{\gamma\in\mathscr{PC}_{1}(\Sigma_{1,1})\mid i(\gamma)=1,\,\ell_{\omega}(\gamma)=5\}|=8$.
@@
It is straightforward from the proposition that $P_{1}(L)=2\cdot|\{$aperiodic necklaces of positive integers $n_{1},\ldots,n_{k}]$ with small variation such that $k+\sum_{i=1}^{k}n_{i}=L-4\}|$. Hence, from the proof of Theorem I.1, that is
@@
Without loss of generality, we assume throughout the proof that $x\leq y$. Let $k\in\mathbb{Z}_{\geq 1}$, and let $n_{i}]_{i}=[n_{1},\dots,n_{k}]$ be a necklace such that $n_{i}\in\{m,m+1\}$ for all $1\leq i\leq k$, $|\{i\mid n_{i}=m\}|=x$, and $|\{i\mid n_{i}=m+1\}|=y$.
First, we prove that if $x\mid y$ then the necklace $n_{i}]_{i}$ does not have variation $2$. If every run of $m+1$ has size $y/x$, then $n_{i}]_{i}$ has small variation. If there exist two runs of $m+1$ with sizes differing by at least $3$ (for example, y/x1y/x-1 and $y/x+2$), then $n_{i}]_{i}$ does not have variation $2$. Thus, if $n_{i}]_{i}$ has variation $2$, then the sizes of its runs of $m+1$ can only take values in $\{y/x-1,y/x,y/x+1\}$, as the sum of the runs have to sum $y$ and there are $x$ of them.
Further, if there is one run of size $y/x+1$, then there is at least one run of size y/x1y/x-1 and there can only be one as otherwise, these give immediately two essential pairs of blocks. Therefore, for a necklace with $2$-variation, there is one run of size y/x1y/x-1, one of size $y/x+1$, and x2x-2 runs of size $y/x$.
If $x>2$, then the necklace cannot have $2$-variation, as one essential pair of blocks is given straight by the runs with difference $2$, i.e. $\{m,m+1,\overset{y/x-1}{\dots},m+1,m\}$ and $\{m+1,\overset{y/x+1}{\dots},m+1\}$, and another one by extending these sets to an adjacent gap of size $y/x$ (e.g. see Figure 7(a)). When $x=2$, there are only two runs, of sizes y/x1y/x-1 and $y/x+1$, and one unique possible configuration with these, giving always variation $2$ (see Figure 7(b)).
@@
Moreover, the necklace $n_{i}]_{i}$ has $2$-variation if and only if the associated run necklace $A[n_{i}]_{i}$ described in Figure 2(c) has. This comes naturally from the map $A$ as two blocks of the form $S_{1}=\{m,n_{j+2},\ldots,n_{j+s-1},m\}$ and $S_{2}=\{m+1,n_{j+2},\ldots,n_{j+s-1},m+1\}$ will map to the same sequence of runs with the first and last one being bigger by one at the second case. Conversely, if there is such a sequence of runs, finding the associated pair of blocks in $n_{i}]_{i}$ is straightforward.
@@
Now, note that since $S_{1}(0)=1$, $|S_{1}(L)|$ can be uniquely determined by (III.4) and (III.5) for any $L\in\mathbb{Z}_{\geq 1}$. On the other hand, a direct computation shows that the function defined by $L\mapsto\lceil\varphi(L)/2\rceil-1$ maps 0 to $1$, and satisfies (III.4) and (III.5). Hence, for any $L\in\mathbb{Z}_{\geq 1}$, we have
@@
$=\begin{cases}8\,\varphi(L-4)&\text{for odd }L,\\ 8\left(\varphi(L-4)+1+\left\lceil\varphi(L/2)/2\right\rceil-1\right)&\text{for even }L.\end{cases}$
@@ IV.  All curves
This arises from the following reasoning. We assume $x_{1}=a$. The exponent $n_{1}$ can be any positive integer, giving a factor $t+t^{2}+\cdots=t/(1-t)$. Next, $n_{2}$ can also be any positive integer, and $x_{2}$ can be chosen between $b$ and $b^{-1}$. This gives a factor 2t/(1t)2t/(1-t), and so on so forth. Finally, $x_{1}$ can also be $a^{-1}$, $b$, or $b^{-1}$, which gives a factor of $4$.
@@
where the term 1/(1t)1/(1-t) corresponds to the case $k=1$. This can be rewritten as
@@
and therefore, as primitive words of length $n$ and primitive words of length $n$ are in $n$-to-11 correspondence, the number of primitive reduced necklaces in $\{a,b,a^{-1},b^{-1}\}$ is
@@
[…]